OpenStudy (anonymous):

What is the solution of the equation? 9=(x-1)^2/3

OpenStudy (anonymous):

power both side by 3

OpenStudy (anonymous):

\[(x-1)^{\frac{2}{3}} = \sqrt[3]{(x-1)^2}\]

OpenStudy (anonymous):

then take sqrt of both side

OpenStudy (anonymous):

so \( 9^3 = (x-1)^2 \)

OpenStudy (anonymous):

then add +1 for both side

OpenStudy (anonymous):

then you will have x=.....

OpenStudy (anonymous):

26?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

help

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

this is an sos